# A Survey on Tabular Data Generation: Utility, Alignment, Fidelity, Privacy, and Beyond

### v1

@@ Abstract
Generative modelling has become the standard approach for synthesising tabular data. However, different use cases demand synthetic data to comply with different requirements to be useful in practice. In this survey, we review deep generative modelling approaches for tabular data from the perspective of four types of requirements: utility of the synthetic data, alignment of the synthetic data with domain-specific knowledge, statistical fidelity of the synthetic data distribution compared to the real data distribution, and privacy-preserving capabilities. We group the approaches along two levels of granularity: (i) based on the primary type of requirements they address and (ii) according to the underlying model they utilise. Additionally, we summarise the appropriate evaluation methods for each requirement and the specific characteristics of each model type. Finally, we discuss future directions for the field, along with opportunities to improve the current evaluation methods. Overall, this survey can be seen as a user guide to tabular data generation: helping readers navigate available models and evaluation methods to find those best suited to their needs.
@@
CWGAN [?] Utility GAN 36
OCT-GAN [?] GAN 59 medium
GANBLR [?] GAN & BN 55
TabGFN [?] GAN & Flow Networks 37
TabDDPM [?] Diffusion 51 slow
STaSy [?] Diffusion 58 slow
CoDi [?] Diffusion 31 medium
GReaT [?] LLM 47 slow
TabPFGen [?] Energy-based model 77
ITS-GAN [?] Alignment GAN & AE 45
C3-TGAN [?] GAN & BN 54
GOGGLE [?] VAE & GNN 168 medium
TDGGD [?] Diffusion 44
C-DGM [?] Neural Network Layer 109 fast ✓✓
DGM+DRL [?] Neural Network Layer 64 fast ✓✓
TGAN [?] Fidelity GAN 55 medium
CTGAN [?] GAN 785 fast
CTAB-GAN [?] GAN 55 medium
CTAB-GAN+ [?] GAN 55 medium
TVAE [?] VAE 785 fast
Neural Spline Flows [?] Normalising Flows 50
FinDiff [?] Diffusion 84 slow
AutoDiff [?] Diffusion & AE 61 slow
TabSyn [?] Diffusion & VAE 48 medium
medGAN [?] Privacy GAN & AE 1071 fast
IT-GAN [?] GAN & AE 59 slow
TableGAN [?] GAN 32 medium
PATE-GAN [?] GAN 617 fast
CuTS [?] GAN 20
Table 1: Overview of the surveyed methods proposed for synthesising tabular data. For each work, we report (i) the requirement type it primarily focuses on, (ii) the model type, (iii) whether the model allows for mixed data types (marked with “✓” if so, and with “✗” otherwise), (iv) the maximum number of features used for evaluation (indicating whether the model can handle high-dimensionality data), (v) the model’s data generation time relative to the other models. In the last four columns, we indicate with “✓” (resp., “✗”) the requirements that were (resp., were not) addressed, specifically, (vi) utility, (vii) alignment, (viii) fidelity, and (ix) privacy. Note that “—” indicates that data are not available, and “✓✓” indicates that the method provides guarantees that the domain-specific constraints are satisfied.
1 Introduction
In recent years, the synthesis of realistic tabular data has emerged as a critical research area, driven by the need for privacy-preserving machine learning, robust AI benchmarking, and data augmentation in high-stakes applications such as finance and healthcare. These diverse needs led to not only a proliferation of different generative models, but also to an explosion of diverse evaluation metrics whose aim is to evaluate whether a model is able to address all the requirements of the given use case. However, the sheer diversity of these metrics has made it increasingly difficult to assess and compare different synthesisers in a standardised manner.
This survey takes a unique stance and analyses models and evaluations methods alike from the point of view of the user, categorising them on the ground of the requirements they primarily address and measure: utility, alignment, fidelity, and privacy. The analysis thus serves as a structured resource for both newcomers and experienced practitioners seeking a formal evaluation framework for tabular data synthesis. To this end, we first give an overview of the tabular data synthesis problem and its requirements (§2). Then, we examine in detail existing methods along two dimensions, grouping them (i) by the requirement they primarily focus on and (ii) by the model architecture type they are based on (§3-6). For each requirement, we detail the relevant evaluation protocols and metrics. A summary of the approaches surveyed in these is given in Table 1, which also gives an overview of the characteristics for each model, such as their ability to handle high-dimensional and mixed data types, as well as their sample generation times. We lastly review other deep generative models originally developed for other fields (such as natural language processing and computer vision), which can be adapted to tabular data synthesis (§7). We conclude with a discussion on related surveys and future directions and opportunities for improvement (§8).
2 Synthesising Tabular Data
Tabular Dataset.
In the tabular data synthesis field, a tabular dataset $\mathcal{D}$ can be usually defined as a triple $\mathcal{D}=(\mathcal{\mathcal{X},\mathcal{A},\mathcal{V}})$ where: (i) $\mathcal{X}=\{{x}^{1},\ldots,{x}^{N}\}$ is a finite set of instances (or rows), with each $\mathbf{x}^{i}$ (with $i=1,\ldots,N$) representing an individual data point, (ii) $\mathcal{A}=\{A^{1},\ldots,A^{K}\}$ is a finite set of attributes (or columns), each representing a feature of the data points, and (iii) $\mathcal{V}:\mathcal{X}\times\mathcal{A}\to\mathbb{D}$ is a value assignment function, which maps every data-point-feature pair $({x}^{i},A^{j})$ into an element of the domain $\mathbb{D}^{j}$, i.e., the set of permissible values for the attribute $A^{j}$. One of the many challenges of this field is given by the fact that the domain of each feature can be very different from one another. Common domains are: binary domains, i.e., $\mathbb{D}=\{0,1\}$, categorical domains, e.g., $\mathbb{D}=\{\text{Red},\text{Yellow},\text{Blue}\}$, and numerical domains, i.e., $\mathbb{D}\subseteq\mathbb{R}$. Given a dataset $\mathcal{D}$ with $N>1$, it is always possible to partition its instances in order to obtain a training set used to train a machine learning model, and a test set used to test a machine learning model.
@@ 2 Synthesising Tabular Data
Deep Generative Modelling for Tabular Data. To formulate the problem of tabular data generation, the literature makes the assumptions that given a tabular dataset $\mathcal{D}$, there exists an unknown distribution $p_{X}$ over the random variables $X\in\bigcup_{j=1}^{K}\mathbb{D}^{j}$ from which $\mathcal{D}$ was created by drawing $N$ i.i.d. samples. The goal of standard generative modeling is to learn from $\mathcal{D}$ the parameters $\theta$ of a generative model such that the model distribution $p_{\theta}$ approximates $p_{X}$. While synthesising tabular data is a long-standing problem [?], in this survey, we focus on methods that use deep generative models, i.e., models based on deep neural networks.
@@
1. Utility requirement: Synthetic data should yield similar predictive performance to real data when used to train machine learning (ML) models for the same task, such as classification or regression. For instance, given a healthcare dataset like WiDS [?], which contains medical records of patients from the first 24 hours of intensive care, and the target column specifies whether a patient has been diagnosed with diabetes mellitus, training a classifier on synthetic data should yield comparable (or better) performance on the real test set to that obtained by training the classifier on real data of the same size. Hence, given an ML model $m$ and a predictive performance metric $r$ (e.g., accuracy), maximising utility entails learning from $\mathcal{D}$ the parameters $\theta$ of a generative model such that training $m$ on the dataset $\mathcal{D}^{\prime}$ sampled from $p_{\theta}$ leads to the best score according to $r$.
2. Alignment requirement: Synthetic data should align with any known (user-provided) domain-specific knowledge. The knowledge can be expressed in multiple ways: simply as constraints on the range of values that a single feature can assume, or as more complex constraints capturing relations between the features. Continuing the example above, if the minimum and the maximum haemoglobin level are two of the features recorded for each patient, then clearly the real data do not contain any records for which the value of the minimum is higher than the value of the maximum level. This type of constraint is easily captured using a linear inequality. Hence, given a set of constraints expressing some background knowledge about the sample space of $p_{X}$, i.e., stating which samples are admissible and which are not, the goal is to learn the parameters $\theta$ of a generative model such that (i) the model distribution $p_{\theta}$ approximates $p_{X}$, and (ii) the sample space of $p_{\theta}$ is compliant with the constraints.

### v2

# A Survey on Deep Learning Approaches for Tabular Data Generation: Utility, Alignment, Fidelity, Privacy, Diversity, and Beyond

@@ Abstract
Generative modelling has become the standard approach for synthesising tabular data. However, different use cases demand synthetic data to comply with different requirements to be useful in practice. In this survey, we review deep generative modelling approaches for tabular data from the perspective of five types of requirements: utility of the synthetic data, alignment of the synthetic data with domain-specific knowledge, statistical fidelity of the synthetic data distribution compared to the real data distribution, privacy-preserving capabilities, and sampling diversity. We group the approaches along two levels of granularity: (i) based on the requirements they address and (ii) according to the underlying model they utilise. Additionally, we summarise the appropriate evaluation methods for each requirement, the relationships among the requirements, and the specific characteristics of each model type. Finally, we discuss future directions for the field, along with opportunities to improve the current evaluation methods. Overall, this survey can be seen as a user guide to tabular data generation: helping readers navigate available models and evaluation methods to find those best suited to their needs.
Start: Select GeneratorSensitive data?High-dimensionality? (>55>55 features)FastInference? (< 1s per 10k samples)FastInference? (< 1s per 10k samples)MixedTypes?MixedTypes?MixedTypes?MixedTypes?Kernel NN:• DP-MERF• DP-NTKGAN:PATE-GAN GAN & AE:medGANIterative NN:GEM+ GAN & AE:IT-GANDiffusion: FinDiff Transf. & GAN: TabTransGAN GAN:• TableGAN• CTAB-GAN+ Iterative NN:GEMGAN: CuTS Diffusion: TDGGD Need constraints? High-dimensionality? (>55>55 features)Fast inference?(< 1s per 10k samples)Fast inference?(< 1s per 10k samples)Mixed Types?Mixed Types?Mixed Types?Mixed Types?PrioritiseDiversity?PrioritiseDiversity?PrioritiseDiversity?PrioritiseDiversity?Energy: TabPFGen VAE & GNN: GOGGLE GAN: • CTGAN • OCT-GAN VAE: TVAE Diffusion: STaSy Diff. & AE: AutoDiffTransf.: TabPFN GAN & BN: GANBLRFlows: Neural Spline Flows Diffusion: CoDi Diff. & VAE: TabSyn GAN: • CWGAN • TGAN • CTAB-GAN GAN & BN: C3-TGAN GAN & AE: ITS-GAN Diffusion: TabDDPM LLM: • GReaT • TabuLa Need constraints? Constraints-enforcing Layer:
• C-DGM
• DGM+DRL
YesYesNoYesNoYesNoYesNoYesYesNoNoNoYesNoYesNoYesNoYesNoYesYesNoYesYesNoYesNoYesNoYesNo
Figure 1: Simplified roadmap to help select an appropriate tabular data synthesiser based on specific requirements, as well as other considerations such as whether the data is high-dimensional and whether the sampling time needs to be low. Note that in this roadmap, we consider fast the models that take less than 1 second to generate 10k samples.
1 Introduction
In recent years, the synthesis of realistic tabular data has emerged as a critical research area, driven by the need for privacy-preserving machine learning, robust AI benchmarking, and data augmentation in high-stakes applications such as finance and healthcare. The successful deployment of deep generative modelling led to synthesisers with the ability to improve utility in downstream tasks (e.g., [Kim et al., 2023]), reduce data scarcity by augmenting existing datasets (e.g., [Jia et al., 2024]), and create new datasets in scenarios where privacy is an important concern (e.g., [Vero et al., 2024]) and anonymising data is a challenging task.
Yet, synthesising such data comes with unique challenges that distinguish it from unstructured domains like computer vision or natural language. Indeed, differently from images or text, where data is homogeneous and presents spatial or sequential correlations, tabular data is inherently heterogeneous, comprising a complex mix of continuous, discrete, and categorical features with diverse distributions. Further, the relations between these features are often non-linear and lack the local structures and patterns leveraged by standard deep learning architectures such as convolutional neural networks. Hence, synthesisers for tabular data face the complex task of faithfully reconstructing these correlations while also handling inherent irregularities such as data imbalance. Moreover, once the data is generated, evaluating its quality proves challenging [Naeem et al., 2020; Räisä et al., 2025], particularly regarding whether statistical inferences can be faithfully reproduced [Montoya Perez et al., 2024], as there is no single metric capable of capturing all the desired requirements.
These diverse needs and technical challenges have led to not only a proliferation of different generative models, but also to an explosion of diverse evaluation metrics whose aim is to evaluate whether a model is able to address all the requirements of the given use case. However, the sheer diversity of these metrics has made it increasingly difficult to assess and compare different synthesisers in a standardised manner.
This survey analyses models and evaluations methods alike from the point of view of the user, categorising them on the ground of the requirements they address and measure: utility, alignment, fidelity, privacy, and diversity. The analysis thus serves as a structured resource for both newcomers and experienced practitioners seeking a formal evaluation framework for tabular data synthesis. To this end, we first give an overview of the tabular data synthesis problem and its requirements (§2). Then, we detail each requirement along with its evaluation protocols and metrics (§3), but also the relationships among these requirements (§3.6). Next, we survey the existing deep learning generative models for synthesising tabular data (§4), grouping them by their model architecture, but also by the requirements they address, as shown in Table 1 and in Figure 3. In particular, Table 1 can be seen as a summary of the surveyed approaches (ordered by their date of publication), providing also an overview of the characteristics for each model, such as their ability to handle high-dimensional and mixed data types, as well as their sample generation times. We lastly review other deep generative models originally developed for other fields (such as natural language processing and computer vision), which can be adapted to tabular data synthesis (§4.6). We conclude with a discussion on related surveys and future directions and opportunities for improvement (§5).
Crucially, throughout this survey, we analyse the key requirements that practitioners typically face when synthesising data. Building on this analysis, in Figure 1, we present a roadmap designed to guide readers in selecting the appropriate synthesiser for their specific use case. This roadmap thus serves as a simplified visualisation of the surveyed methods, with methods’ grouping based on critical considerations such as data sensitivity, high-dimensionality, sampling speed, but also the need for the generated data to adhere to domain constraints. Notably, the decision to enforce domain constraints is positioned as a final step. Indeed, in case constraints are needed, the listed methods to guarantee their satisfaction can be simply built on top of a given synthesiser as they are differentiable layers able to be integrated into any neural network model during training. Conversely, the decision regarding privacy is fundamental and, thus, appears as the first step in the roadmap, specifying whether a specialised privacy-preserving mechanism is required from the start.
2 Synthesising Tabular Data
Tabular Dataset.
In the tabular data synthesis field, a tabular dataset $\mathcal{D}$ is a collection of $N$ i.i.d. samples (indexed by $i\in\{1,\ldots,N\}$ such that the $i$-th sample constitutes the $i$-th row) drawn from an unknown joint distribution $p_{\mathcal{X}}$ over the variables (i.e., columns) $(X_{1},X_{2},\ldots,X_{K})$. Thus, for each $i\in\{1,\ldots,N\}$, the sample $\mathbf{{x}}^{i}=(x_{1}^{i},x_{2}^{i},\ldots,x_{K}^{i})$ is an assignment of values to the variables $(X_{1},X_{2},\ldots,X_{K})$, i.e., $x^{i}_{j}\in\mathbb{D}^{j}$ is the value assigned to variable $X_{j}$, for each $j\in\{1,\ldots,K\}$, where $\mathbb{D}^{j}$ is the set of permissible values for variable $X_{j}$. One of the many challenges of this field is given by the fact that the domain of each feature can be very different from one another. Common domains are: binary domains, i.e., $\mathbb{D}=\{0,1\}$, categorical domains, e.g., $\mathbb{D}=\{\text{Red},\text{Yellow},\text{Blue}\}$, and numerical domains, i.e., $\mathbb{D}\subseteq\mathbb{R}$. Given a dataset $\mathcal{D}$ with $N>1$, it is always possible to partition its instances in order to obtain a training set used to train a machine learning model, and a test set used to test a machine learning model.
Table 1: Overview of the surveyed methods proposed for synthesising tabular data. For each work, we report (i) the model type, (ii) whether the model allows for mixed data types (marked with “✓” if so, and with “✗” otherwise), (iii) the maximum number of features used for evaluation (indicating whether the model can handle high-dimensionality data), (iv) the model’s data generation time relative to the other models. In the last five columns, we indicate with “✓” (resp., “✗”) the requirements that were (resp., were not) addressed, and with “•” the requirements whose resolution depends on the base DGM model upon which the Constraints-enforcing methods are built. Specifically, these requirements are: (v) utility, (vi) background knowledge alignment, (vii) statistical fidelity, (viii) privacy, and (ix) diversity. Regarding the generation time, the three indicators are assigned based on the time (in seconds) it takes to generate 10K samples, as follows: fast for generation time less than 0.5 s, medium for generation time higher than 0.5 s but less than 1 s, and slow, otherwise. Further, note that “—” indicates that data are not available, and “✓✓” indicates that the method provides guarantees that the domain-specific constraints are satisfied. Note that the generation time values have been collected from the surveyed papers, aggregated and had thresholds applied to them to assign them to one of the three possible indicators provided, i.e., fast, medium, and slow.
@@ 2 Synthesising Tabular Data
Diversity
medGAN [Choi et al., 2017] GAN & AE 1071 fast
TGAN [Xu & Veeramachaneni, 2018] GAN 55 medium
TableGAN [Park et al., 2018] GAN 32 medium
PATE-GAN [Yoon et al., 2019] GAN 617 fast
ITS-GAN [Chen et al., 2019] GAN & AE 45
CTGAN [Xu et al., 2019] GAN 785 fast
TVAE [Xu et al., 2019] VAE 785 fast
Neural Spline Flows [Durkan et al., 2019] Normalising Flows 50
CWGAN [Engelmann & Lessmann, 2021] GAN 36
OCT-GAN [Kim et al., 2021] GAN 59 medium
GANBLR [Zhang et al., 2021] GAN & BN 55
CTAB-GAN [Zhao et al., 2021] GAN 55 medium
IT-GAN [Lee et al., 2021] GAN & AE 59 slow
GEM [Liu et al., 2021] Iterative NN 14 fast
DP-MERF [Harder et al., 2021] Kernel-based NN 617 fast
GOGGLE [Liu et al., 2022] VAE & GNN 168 medium
CTAB-GAN+ [Zhao et al., 2022] GAN 55 medium
TabDDPM [Kotelnikov et al., 2023] Diffusion 51 slow
STaSy [Kim et al., 2023] Diffusion 58 slow
CoDi [Lee et al., 2023] Diffusion 31 medium
GReaT [Borisov et al., 2023] LLM 47 slow
TabPFGen [Ma et al., 2023] Energy-based model 77
C3-TGAN [Han et al., 2023] GAN & BN 54
FinDiff [Sattarov et al., 2023] Diffusion 84 slow
AutoDiff [Suh et al., 2023] Diffusion & AE 61 slow
DP-NTK Yang et al. [2024] Kernel-based NN 617 fast
TDGGD [Jia et al., 2024] Diffusion 44
C-DGM [Stoian et al., 2024] Constraints-enforcing 109 fast ✓✓ $\bullet$ $\bullet$ $\bullet$
TabSyn [Zhang et al., 2024] Diffusion & VAE 48 medium
CuTS [Vero et al., 2024] GAN 20
TabuLa [Zhao et al., 2025] LLM 55 slow
DGM+DRL [Stoian & Giunchiglia, 2025] Constraints-enforcing 64 fast ✓✓ $\bullet$ $\bullet$ $\bullet$
TabTransGAN [Zhang et al., 2025] Transformer & GAN 59
TabPFN [Hollmann et al., 2025] Transformer 500 slow
GEM+ [Maddock et al., 2025] Iterative NN 123 fast
Deep Generative Modelling for Tabular Data.
@@
To formulate the problem of tabular data generation, the literature makes the assumptions that given a tabular dataset $\mathcal{D}$, there exists an unknown joint distribution $p_{X}$ over the random variables $(X_{1},X_{2},\ldots,X_{K})$, where $X_{j}\in\mathbb{D}^{j}$, such that $\mathcal{D}$ was created by drawing $N$ i.i.d. samples. The goal of standard generative modelling is to learn from $\mathcal{D}$ the parameters $\theta$ of a generative model such that the model distribution $p_{\theta}$ approximates $p_{X}$.
While synthesising tabular data is a long-standing problem [Reiter, 2005], in this survey, we focus on methods that use deep generative models, i.e., models based on deep neural networks.
@@
1. Utility requirement: Synthetic data should yield similar predictive performance to real data when used to train machine learning (ML) models for the same task, such as classification or regression. For instance, given a healthcare dataset like WiDS [Matthys et al., 2021], which contains medical records of patients from the first 24 hours of intensive care, and the target column specifies whether a patient has been diagnosed with diabetes mellitus, training a classifier on synthetic data should yield comparable (or better) performance on the real test set to that obtained by training the classifier on real data of the same size. Hence, given an ML model $m$ and a predictive performance metric $r$ (e.g., accuracy), maximising utility entails learning from $\mathcal{D}$ the parameters $\theta$ of a generative model such that training $m$ on the dataset $\mathcal{D}^{\prime}$ sampled from $p_{\theta}$ leads to the best score according to $r$.
In addition to predictive performance, utility can be broadly defined as the preservation of valid statistical inferences [Raghunathan et al., 2003]. For instance, in differentially private settings, Rosenblatt et al. [2023] quantify this as the likelihood that the empirical conclusions derived from the real data would not change if synthetic data were used instead. In these cases, maximising this statistical utility boils down to maximising this likelihood.
2. Alignment requirement: Synthetic data should align [^1] To further clarify our categorisation, we acknowledge the possible terminology overlap with previous papers. For example, Xu et al. [2019] mention that the discriminator of CTGAN is trained such that the synthetic data aligns with the real data. However, in this case, the term align is used informally and refers to the process of matching the real data distribution, which falls under the fidelity requirement in our survey. The formal definition of alignment has been introduced in [Stoian et al., 2024], and is reproduced here to convey alignment with the given background knowledge. with any known (user-provided) domain-specific knowledge. The knowledge can be expressed in multiple ways: simply as constraints on the range of values that a single feature can assume, or as more complex constraints capturing relations between the features. Continuing the example above, if the minimum and the maximum haemoglobin level are two of the features recorded for each patient, then clearly the real data do not contain any records for which the value of the minimum is higher than the value of the maximum level. This type of constraint is easily captured using a linear inequality. Hence, given a set of constraints expressing some background knowledge about the sample space of $p_{X}$, i.e., stating which samples are admissible and which are not, the goal is to learn the parameters $\theta$ of a generative model such that (i) the model distribution $p_{\theta}$ approximates $p_{X}$, and (ii) the sample space of $p_{\theta}$ is compliant with the constraints.